OpenStudy (anonymous):

5x+7x/3*5=? when x is 9 i'll give medal and fan

OpenStudy (anonymous):

you have to remember pemdas so what you're gonna do is input the 9 in first 5(9) +7(9) / 3*5 then you do multiplication 45 + 63/ 15 then you do division so 63/15 would be 4.2 then you add 45 and 4.2 and its 49.2

OpenStudy (anonymous):

5⋅9+793⋅5 i thimk not sure thogh isolved by factoring

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

hold on its 5*9+7 9/3*5

OpenStudy (anonymous):

so its 95 is the answer

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no no no.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its not 95 i gave a step by step in my comment.

OpenStudy (anonymous):

its 150 i figured it out

OpenStudy (anonymous):

okay explain it to me then, did you go through pemdas?

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

look it is like 5(9)+7(9)/3*5. first do the devide- now it is 45+7(3)*5 = 45+105=150.

OpenStudy (anonymous):

is it correct @TriniEmo ?

OpenStudy (anonymous):

no you don't divide first pemdas says specifically to multiply before dividing.

OpenStudy (anonymous):

parenthesis exponents multiplication division addition subtraction

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@timo86m @hartnn Please help clear this up.

OpenStudy (anonymous):

Parenthesis Exponents Multiplication Divison Addition Subtraction

hartnn (hartnn):

5x+7x/3*5 when x=9 so we put x=9 5(9) + 7*9/3 *5 45 +7*3*5 45+105 150 is correct :)

OpenStudy (anonymous):

thanks @hartnn

OpenStudy (anonymous):

hartnn i want you to explain how you divide before you multiply

OpenStudy (anonymous):

I always learned to do all multiplication before you do division

hartnn (hartnn):

yes, multiplication and division have equal priority, so when they are together, we go from left to right 7*9/3 = 63/3 which is same as 7*3

OpenStudy (anonymous):

equal priority, really e.e dangit

OpenStudy (anonymous):

well if they have equal priority then it is 150 and i apologize for complications